OpenStudy (anonymous):

how would i solve (7/8)-(1/x)=3/4 ? i got x=8 but i know for a fact thats not right. can someone please show me?

OpenStudy (campbell_st):

it becomes -1/x = -1/8 (subtract 7/8 from both sides) equating denominators x = 8

OpenStudy (anonymous):

so i am right?

OpenStudy (anonymous):

yeah u are right honey :) got the same nswer

OpenStudy (anonymous):

lol thanks. i must have done the checking stage wrong then

OpenStudy (anonymous):

anytime :) just be a little bit careful,is math

OpenStudy (phi):

Check \[\frac{7}{8}-\frac{1}{8}= \frac{6}{8}= \frac{3}{4}\]